# TexGaussian: Generating High-quality PBR Material via Octree-based 3D Gaussian Splatting

### v1

@@ Abstract
Physically Based Rendering (PBR) materials play a crucial role in modern graphics, enabling photorealistic rendering across diverse environment maps. Developing an effective and efficient algorithm that is capable of automatically generating high-quality PBR materials rather than RGB texture for 3D meshes can significantly streamline the 3D content creation. Most existing methods leverage pre-trained 2D diffusion models for multi-view image synthesis, which often leads to severe inconsistency between the generated textures and input 3D meshes. This paper presents TexGaussian, a novel method that uses octant-aligned 3D Gaussian Splatting for rapid PBR material generation. Specifically, we place each 3D Gaussian on the finest leaf node of the octree built from the input 3D mesh to render the multi-view images not only for the albedo map but also for roughness and metallic. Moreover, our model is trained in a regression manner instead of diffusion denoising, capable of generating the PBR material for a 3D mesh in a single feed-forward process. Extensive experiments on publicly available benchmarks demonstrate that our method synthesizes more visually pleasing PBR materials and runs faster than previous methods in both unconditional and text-conditional scenarios, which exhibit better consistency with the given geometry. Our code and trained models are available at https://3d-aigc.github.io/TexGaussian.
@@ 1 Introduction
  • We propose TexGaussian, a novel PBR material generation method based on the above two techniques. To our knowledge, our method first generates PBR material directly in 3D space. Qualitative and quantitative experiments have been conducted to verify the superiority of the quality and efficiency of our method over other existing approaches.
@@ 2.1 Multi-view Images Synthesis
Many previous works have tried to leverage the powerful T2I model to assist texture generation for 3D shapes. Specifically, they render the depth map of input 3D mesh from multiple views and use depth conditional T2I models [56] to synthesize RGB images and perform text-conditioned texture synthesis. TEXTure [40] and Text2Tex [7] iteratively paint a mesh from different views. However, images synthesized in early view could produce errors that are not reconcilable with the geometry that is observed in later views. Many subsequent works try to alleviate multi-view inconsistency via different alignment modules. TexFusion [5] proposes a sequential interlaced multi-view sampler that interleaves texture assembling with denoising steps in different camera views. Similarly, TexGen [19] directly enforces view consistent sampling in RGB texture space and develops a noise resampling strategy to retain rich texture details. TexPainter [55] blends images from different views into a common color-space texture image by weighted averaging to guarantee multi-view consistency. GenesisTex [15] introducing style consistency and dynamic alignment across multiple viewpoints. To remove light influence from 2D diffusion models, Paint3D [54] contribute separate UV Inpainting and UVHD diffusion models specialized in shape-aware refinement. Although these methods achieve impressive texture results, they can still hardly comprehend the overall geometry of input 3D mesh.
@@ 2.2 Optimization-based 3D Generation
Before the emergence of large-scale Text-to-Image generative model, earlier methods [18, 9, 29, 30, 32] propose to optimize texture map of 3D object via natural language supervised visual model, CLIP [38]. Subsequently, score distillation sampling (SDS) was adopted by DreamFusion [37] and Magic3D [24]. The key idea is to optimize 3D representations such as NeRF [31] or InstantNGP [33] with the gradient guidance from 2D diffusion priors [37, 50]. To generate PBR material, TextureDreamer [52] optimizes spatially-varying bidirectional reflectance distribution (BRDF) field through personalized geometric-aware score distillation. Fantasia3D [8] uses a single predefined environmental. However, the generated images from diffusion models may not be consistent with the given environment light. DreamMat [58] proposes a novel geometry and light-aware diffusion model, which is trained to generate images that are consistent with the given environment light. FlashTex[13] also proposes a light-conditioned diffusion model within a two-stage pipeline, combining reconstruction and SDS optimization to enhance texture quality and achieve better light disentanglement. However, these methods struggle with the Janus problem due to the semantically ambiguous. And the time consumption is relatively too long to use in practice.
@@ 3 Method
In this section, we provide a detailed explanation of our proposed method, TexGaussian. The overall pipeline of our method is shown in Fig 1. Existing texture synthesis approaches mainly rely on pre-trained 2D diffusion models, which struggle to fully understand the overall 3D structure. This often leads to misalignment between the generated texture map and 3D semantics. Our goal is to synthesize high-quality PBR materials for a given mesh directly in 3D space.
@@ 3.2 Octant-aligned 3D Gaussian Splatting
We pre-fitting the parameters of 3D Gaussians on the constructed octree for each 3D object in our dataset via the original loss function in [21] on multi-view RGB images, roughness maps, and metallic maps:
@@ 3.3 Octree-based 3D U-Net Training
To effectively train our 3D U-Net, we adopt the regressive loss objective, which could further facilitate the generation process. The input to our octree-based 3D U-Net is the geometry feature on each octree’s finest leaf node, such as normal and local displacement. For text-conditioned PBR material synthesis, the text feature is extracted by pre-trained CLIP model [38] and is fed to U-Net via the octree-based multi-head cross attention mechanism similar to [47]. The predicted 3D Gaussians are rasterized from multiple views via Gaussian Splatting [21]. At each training step, we rasterize the RGB images, alpha images, roughness and metallic maps from randomly selected eight views. Following [46], we apply mean square error (MSE) loss and VGG-based LPIPS loss [57] to the RGB image, roughness map, and metallic map:
$L_{\text{RGB}}=L_{\text{MSE}}(I_{\text{RGB}},I_{\text{RGB}}^{\text{GT}})+L_{\text{LPIPS}}(I_{\text{RGB}},I_{\text{RGB}}^{\text{GT}}),$ (2)
@@
$L_{\text{total}}=L_{\text{RGB}}+L_{\text{R}}+L_{\text{M}}+L_{\alpha}+L_{\text{3D}}.$ (6)
@@ 3.4 Texture and Material Baking
In the inference stage, we also first build the corresponding octree for the input 3D mesh. Then, we use our trained octree-based 3D U-Net to generate 3D Gaussian on every octant and rasterize them from multiple views. The ultimate output of our method should be a global texture map and material map. Thus, we rasterize the input mesh using the differentiable renderer [23] and optimize its albedo, roughness and metallic parameters via the MSE loss between the Nvdiffrast [23] rendering results and 3D Gaussian rendering results. With adequately optimized implementation, this process takes only about several seconds to bake the multi-view images to untextured 3D model. After the optimization, our input mesh paired with its albedo and material map is capable of performing physically based rendering in new illumination environments.
@@ 4.1 Implementation Details
We train our TexGaussian model on two publicly-available datasets: ShapeNet [6] and Objaverse [12]. For 3D objects in the ShapeNet dataset, since they only contain albedo maps without PBR materials, our model is trained to generate RGB textures only. We train our model on four categories of ShapNet: bench, car, chair and table which is consistent with previous works [53, 25]. We curate a subset of Objaverse models encoded with PBR materials and convert those using the specular-glossiness workflow to the metallic-roughness workflow to achieve a consistent PBR representation. This process resulted in a total of $29,200$ models. For each 3D object, we render its RGBA image, roughness map, and metallic map from 64 views of $512^{2}$ for training.
@@
Table 1: Quantitative comparison of the FID and KID (×$\times 10^{2}$) score as well as the inference time of TexGaussian and other methods on ShapeNet dataset [6]. The top part reports the comparison with methods that also train per-category models on ShapeNet. The bottom part reports the comparison with 2D diffusion-based methods, which select $50$ samples from the test set of each category.
@@ 4.2 Unconditional RGB Texture Generation
Figure 3: Examples of generated RGB texture obtained by TexGaussian and other state-of-the-art models on the same 3D object. Please zoom in for a better inspection.
Figure 4: Qualitative comparison with Fantasia3D [8], FlashTex [13] and DreamMat [58]. We provide the rendered image, albedo map, roughness map, and metallic map for each 3D object.
@@ 4.3 Text-conditioned PBR Material Generation
We train the text-conditioned PBR material generation model on our filtered subset of Objaverse [12]. We use the text descriptions of 3D objects from Cap3D [27, 28] to train our model. We choose $29,000$ 3D objects in our filtered subset for training and the rest $200$ for testing. Fig LABEL:fig:teaser shows some generative results by proposed TexGaussian on the test set. We can see that our model comprehends the overall geometry feature of input 3D mesh. It is capable of synthesizing high-quality albedo, roughness, and metallic, which are aligned well with 3D semantics such as accents with high metallic on the wooden box and rope with high roughness on the jug.
@@
We conduct quantitative comparison with three state-of-the-art text-conditioned PBR material synthesis methods: Fantasia3D [8], FlashTex [13], and DreamMat [58] on our test set. Table 2 reports FID and KID scores on both multi-view albedo map and PBR rendering images under the same illumination environment. TexGaussian outperforms all the baseline methods in achieving the best visual quality of the generated appearances. We also report the average inference time of different methods on a single NVIDIA A100 (40G) GPU across our test set. Due to the iterative optimization of score distillation sampling, all other three methods cost at least 20 minutes for the generation. In contrast, our TexGaussian only takes about 20 seconds, in which one second is for predicting the Gaussian parameters and the rest is for baking, which results in 60×60$\times$60 × faster compared to previous approaches.
@@
Qualitative Comparison
Fig 4 visualizes the generated albedo, roughness, and metallic of each compared method from the same text prompt and untextured meshes. We also show the PBR rendering images of the generated materials under the same environment light. It can be observed that Fantasia3D [8] and FlashTex [13] generate irregular colors on mesh surfaces, which lead to rendering results with low quality. DreamMat [58] is capable of generating a visually pleasing appearance for test 3D models. However, it can hardly align the generated PBR material well with the 3D semantics, such as the black band on the wooden barrel. What’s more, it tends to generate over-saturated colors, which is demonstrated in the wooden chest and green grenade. In addition, all of the compared methods struggle to completely disentangle the light and texture, which results in relatively chaotic results of generated metallic and roughness. On the contrary, our method is directly trained from 3D origin data. It is capable of generating clean and smooth PBR material while fully comprehending the overall 3D structure. We provide more generative results in the supplementary material.
@@ 4.4 Ablation Study
For the purpose of analyzing the impact of different designs in our model, we conduct ablation studies by removing or changing some proposed modules.
@@
$L_{\text{2D}}=L_{\text{RGB}}+L_{\text{R}}+L_{\text{M}}+L_{\alpha}.$ (7)
We train two additional TexGaussian models using only the $L_{\text{2D}}$ or $L_{\text{3D}}$ loss on ShapeNet car category due to its large variations and complexity of texture to validate their effectiveness. It is worth noting that in this dataset, $L_{\text{2D}}=L_{\text{RGB}}+L_{\alpha}$ due to the lack of material information. The training curves of MSE loss and LPIPS loss between rendering images and ground-truth ones are shown in Fig 6, which demonstrate the effects of $L_{2D}$ and $L_{3D}$. From the loss curves, we can conclude that $L_{\text{3D}}$ facilitates the process of convergence and $L_{2D}$ enhances the quality of synthesized images, as evidenced by the large margin of improvement in the LPIPS loss when introducing $L_{2D}$. We also provide some qualitative results to verify the effectiveness of $L_{2D}$ on the test set of car category in Fig 7. Only using $L_{3D}$ results in a relatively blurry texture map due to the discreteness of 3D Gaussian which is similar to the characteristic of 3D point clouds analyzed above.
@@
Figure 7: Results of our model trained with and without $L_{2D}$. The green rectangles highlight the shortcomings of TexGaussian model trained without $L_{2D}$.
@@ Appendix A Preliminary of 3D Gaussian Splatting
$\boldsymbol{\Sigma}=\boldsymbol{RSS}^{\top}\boldsymbol{R}^{\top},$ (9)
@@
$\boldsymbol{\mu}^{\prime}=\boldsymbol{K}[\boldsymbol{\mu},1]^{\top},\quad\boldsymbol{\Sigma}^{\prime}=\boldsymbol{J}\boldsymbol{T}\boldsymbol{\Sigma}\boldsymbol{T}^{\top}\boldsymbol{J}^{\top},$ (10)
@@ Appendix B Network Details
The network architecture of the octree-based 3D U-Net we used in unconditional RGB texture generation is shown in Fig 8. The U-Net has five hierarchical levels, corresponding to octree depths of 8, 7, 6, 5 and 4, with resolutions of $256^{3},128^{3},64^{3},32^{3},16^{3}$. The feature dimensions are set to $32,64,128,256,256$ respectively. The channel of input and output feature is $4$ and $13$ as described in the main manuscript.
@@
The network architecture of the octree-based 3D U-Net we used in text-conditioned PBR material generation is shown in Fig 9. The U-Net has five hierarchical levels, corresponding to octree depths of 8, 7, 6, 5 and 4, with resolutions of $256^{3},128^{3},64^{3},32^{3},16^{3}$. The feature dimensions are set to $64,128,256,512,512$ respectively. The text feature is fed to U-Net via the octree-based multi-head cross attention mechanism. The cross attention layers are only inserted at the least two down-sampling blocks, the middle block and the two first up-sampling blocks to save GPU memory.

### v2

@@ Abstract
Physically Based Rendering (PBR) materials play a crucial role in modern graphics, enabling photorealistic rendering across diverse environment maps. Developing an effective and efficient algorithm that is capable of automatically generating high-quality PBR materials rather than RGB texture for 3D meshes can significantly streamline the 3D content creation. Most existing methods leverage pre-trained 2D diffusion models for multi-view image synthesis, which often leads to severe inconsistency between the generated textures and input 3D meshes. This paper presents TexGaussian, a novel method that uses octant-aligned 3D Gaussian Splatting for rapid PBR material generation. Specifically, we place each 3D Gaussian on the finest leaf node of the octree built from the input 3D mesh to render the multi-view images not only for the albedo map but also for roughness and metallic. Moreover, our model is trained in a regression manner instead of diffusion denoising, capable of generating the PBR material for a 3D mesh in a single feed-forward process. Extensive experiments on publicly available benchmarks demonstrate that our method synthesizes more visually pleasing PBR materials and runs faster than previous methods in both unconditional and text-conditional scenarios, exhibiting better consistency with the given geometry. Our code and trained models are available at https://3d-aigc.github.io/TexGaussian.
@@ 1 Introduction
  • We propose TexGaussian, a novel PBR material generation method based on the above two techniques. To our knowledge, our work is the first to generate PBR materials directly in 3D space. Qualitative and quantitative experiments have been conducted to verify the superiority of the quality and efficiency of our method over other existing approaches.
@@ 2.1 Multi-view Images Synthesis
Many previous works have tried to leverage the powerful T2I model to assist texture generation for 3D shapes. Specifically, they render the depth map of input 3D mesh from multiple views and use depth conditional T2I models [58] to synthesize RGB images and perform text-conditioned texture synthesis. TEXTure [41] and Text2Tex [8] iteratively paint a mesh from different views. However, images synthesized in early view could produce errors that are not reconcilable with the geometry that is observed in later views. Many subsequent works try to alleviate multi-view inconsistency via different alignment modules. TexFusion [6] proposes a sequential interlaced multi-view sampler that interleaves texture assembling with denoising steps in different camera views. Similarly, TexGen [20] directly enforces view consistent sampling in RGB texture space and develops a noise resampling strategy to retain rich texture details. TexPainter [57] blends images from different views into a common color-space texture image by weighted averaging to guarantee multi-view consistency. To remove light influence from 2D diffusion models, Paint3D [56] contribute separate UV Inpainting and UVHD diffusion models specialized in shape-aware refinement. More recently, Meta 3D TextureGen [1] utilizes geometry-aware text-to-image model to generate a multi-view image for texture. Although these methods achieve impressive texture results, they can still hardly comprehend the overall geometry of input 3D mesh.
@@ 2.2 Optimization-based 3D Generation
Before the emergence of large-scale Text-to-Image generative model, earlier methods [19, 10, 30, 31, 33] propose to optimize texture map of 3D object via natural language supervised visual model, CLIP [39]. Subsequently, score distillation sampling (SDS) was adopted by DreamFusion [38] and Magic3D [25]. The key idea is to optimize 3D representations such as NeRF [32] or InstantNGP [34] with the gradient guidance from 2D diffusion priors [38, 51]. To generate PBR material, TextureDreamer [53] optimizes spatially-varying bidirectional reflectance distribution (BRDF) field through personalized geometric-aware score distillation. Fantasia3D [9] uses a single predefined environmental. However, the generated images from diffusion models may not be consistent with the given environment light. Another concurrent work Paint-it [54] is similar to Fantasia3D [9] but with Deep Convolutional Physically-Based Rendering (DC-PBR) parameterization. DreamMat [60] proposes a novel geometry and light-aware diffusion model, which is trained to generate images that are consistent with the given environment light. FlashTex[14] uses a light-conditioned diffusion model within a two-stage pipeline, combining reconstruction and SDS optimization to enhance texture quality and achieve better light disentanglement. However, these methods struggle with the Janus problem due to the semantically ambiguous. And the time consumption is relatively too long to use in practice.
@@ 3 Method
In this section, we provide a detailed explanation of our proposed method, TexGaussian. The overall pipeline of our method is shown in Fig. 1. Existing texture synthesis approaches mainly rely on pre-trained 2D diffusion models, which struggle to fully understand the overall 3D structure. This often leads to misalignment between the generated texture map and 3D semantics. Our goal is to synthesize the high-quality PBR material for a given mesh directly in 3D space.
@@ 3.2 Octant-aligned 3D Gaussian Splatting
We pre-fitting the parameters of 3D Gaussians on the constructed octree for each 3D object in our dataset via the original loss function in [22] on multi-view albedo images, roughness maps, and metallic maps:
@@ 3.3 Octree-based 3D U-Net Training
To effectively train our 3D U-Net, we adopt the regressive loss objective, which could further facilitate the generation process. The input to our octree-based 3D U-Net is the geometry feature on each octree’s finest leaf node, such as normal and local displacement. For text-conditioned PBR material synthesis, the text feature is extracted by pre-trained CLIP model [39] and is fed to U-Net via the octree-based multi-head cross attention mechanism similar to [48]. The predicted 3D Gaussians are rasterized from multiple views via Gaussian Splatting [22]. At each training step, we rasterize the albedo images, alpha images, roughness and metallic maps from randomly selected eight views. Following [47], we apply the mean square error (MSE) loss and the VGG-based LPIPS loss [59] to the albedo image, roughness map, and metallic map:
$L_{\text{albedo}}=L_{\text{MSE}}(I_{\text{albedo}},I_{\text{albedo}}^{\text{GT}})+L_{\text{LPIPS}}(I_{\text{albedo}},I_{\text{albedo}}^{\text{GT}}),$ (2)
@@
$L_{\text{total}}=L_{\text{albedo}}+L_{\text{R}}+L_{\text{M}}+L_{\alpha}+L_{\text{3D}}.$ (6)
@@ 3.4 Texture and Material Baking
In the inference stage, we also first build the corresponding octree for the input 3D mesh. Then, we use our trained octree-based 3D U-Net to generate 3D Gaussian on every octant and rasterize them from multiple views. The ultimate output of our method should be a global texture map and material map. Thus, we rasterize the input mesh using the differentiable renderer [24] and optimize its albedo, roughness and metallic parameters via the MSE loss between the Nvdiffrast [24] rendering results and 3D Gaussian rendering results. With adequately optimized implementation, this process takes only about several seconds to bake the multi-view images to an untextured 3D model. After the optimization, our input mesh paired with its albedo and material map is capable of performing physically based rendering in new illumination environments.
@@ 4.1 Implementation Details
We train our TexGaussian model on two publicly-available datasets: ShapeNet [7] and Objaverse [13]. For 3D objects in the ShapeNet dataset, since they only contain albedo maps without PBR materials, our model is trained to generate RGB textures only. We train our model on four categories of ShapNet: bench, car, chair and table which is consistent with previous works [55, 26]. We curate a subset of Objaverse models encoded with PBR materials and convert those using the specular-glossiness workflow to the metallic-roughness workflow to achieve a consistent PBR representation. This process resulted in a total of $29,200$ models. For each 3D object, we render its RGBA albedo image, roughness map, and metallic map from 64 views of $512^{2}$ for training.
@@ 4.2 Unconditional RGB Texture Generation
Table 1: Quantitative comparison of the FID and KID (×$\times 10^{2}$) score as well as the inference time of TexGaussian and other methods on the ShapeNet dataset [7]. The top part reports the comparison with methods that also train per-category models on ShapeNet. The bottom part reports the comparison with 2D diffusion-based methods, which select $50$ samples from the test set of each category.
@@
Figure 3: Examples of generated RGB textures obtained by TexGaussian and other state-of-the-art models on the same 3D object. Please zoom in for a better inspection.
@@
Figure 4: Qualitative comparison with Fantasia3D [9], FlashTex [14], DreamMat [60] and Make-it-Real [16]. We show the rendered image, albedo map, roughness map, and metallic map for each 3D object.
@@ 4.3 Text-conditioned PBR Material Generation
We train the text-conditioned PBR material generation model on our filtered subset of Objaverse [13]. We use the text descriptions of 3D objects from Cap3D [28, 29] to train our model. We choose $29,000$ 3D objects in our filtered subset for training and the rest $200$ for testing. Fig. LABEL:fig:teaser shows some generative results by proposed TexGaussian on the test set. We can see that our model comprehends the overall geometry feature of input 3D mesh. It is capable of synthesizing high-quality albedo, roughness, and metallic, which are aligned well with 3D semantics such as accents with high metallic on the wooden box and rope with high roughness on the jug.
@@
We conduct quantitative comparison with some state-of-the-art text-conditioned PBR material synthesis methods: Fantasia3D [9], FlashTex [14], and DreamMat [60] on our test set. Table 2 reports FID and KID scores on both multi-view albedo map and PBR rendering images under the same illumination environment. TexGaussian outperforms all the baseline methods in achieving the best visual quality of the generated appearances. We also report the average inference time of different methods on a single NVIDIA A100 (40G) GPU across our test set. Due to the iterative optimization of score distillation sampling, all other three methods cost at least 20 minutes for the generation. In contrast, our TexGaussian only takes about 20 seconds, in which one second is for predicting the Gaussian parameters and the rest is for baking, which performs 60×60$\times$60 × faster than previous methods.
Qualitative Comparison
Fig. 4 visualizes the generated albedo, roughness, and metallic of each compared method from the same text prompt and untextured meshes. We also show the PBR rendering images of the generated materials under the same environment light. It can be observed that Fantasia3D [9] and FlashTex [14] generate irregular colors on mesh surfaces, which lead to rendering results with low quality. DreamMat [60] is capable of generating a visually pleasing appearance for test 3D models. However, it can hardly align the generated PBR material well with the 3D semantics, such as the black band on the wooden barrel. What’s more, it tends to generate over-saturated colors, which is demonstrated in the wooden chest and green grenade. In addition, all of the compared methods struggle to completely disentangle the light and texture, which results in relatively chaotic results of generated metallic and roughness. We also compare our method with a Multimodal Large Language Model based method, Make-it-Real [16], which accepts 3D objects paired with albedo maps generated by our TexGaussian as input. It can be observed that Make-it-Real struggles to capture the overall 3D structure of input meshes. As a consequence, it tends to predict single-value PBR maps, leading to irregular rendering results, such as the wooden chest with extremely large metallic value. On the contrary, our method is directly trained from 3D original data. It is capable of generating clean and smooth PBR material while fully comprehending the overall 3D structure. We provide more generative results in the supplementary material.
@@ 4.4 Ablation Study
$L_{\text{2D}}=L_{\text{albedo}}+L_{\text{R}}+L_{\text{M}}+L_{\alpha}.$ (7)
We train two additional TexGaussian models using only the $L_{\text{2D}}$ or $L_{\text{3D}}$ loss on ShapeNet car category due to its large variations and complexity of texture to validate their effectiveness. It is worth noting that in this dataset, $L_{\text{2D}}=L_{\text{albedo}}+L_{\alpha}$ due to the lack of material information. The training curves of MSE loss and LPIPS loss between rendering images and ground-truth ones are shown in Fig. 6, which demonstrate the effects of $L_{2D}$ and $L_{3D}$. From the loss curves, we can conclude that $L_{\text{3D}}$ facilitates the process of convergence and $L_{2D}$ enhances the quality of synthesized images, as evidenced by the large margin of improvement in the LPIPS loss when introducing $L_{2D}$. We also provide some qualitative results to verify the effectiveness of $L_{2D}$ on the test set of car category in Fig. 7. Only using $L_{3D}$ results in a relatively blurry texture map due to the discreteness of 3D Gaussian which is similar to the characteristic of 3D point clouds analyzed above.
@@
Figure 7: Results of our model trained with and without $L_{2D}$.
@@ Appendix A Preliminary of 3D Gaussian Splatting
$\boldsymbol{\Sigma}=\boldsymbol{RSS}^{\top}\boldsymbol{R}^{\top}.$ (9)
@@
$\boldsymbol{\mu}^{\prime}=\boldsymbol{K}[\boldsymbol{\mu},1]^{\top},\quad\boldsymbol{\Sigma}^{\prime}=\boldsymbol{J}\boldsymbol{T}\boldsymbol{\Sigma}\boldsymbol{T}^{\top}\boldsymbol{J}^{\top}.$ (10)
@@ Appendix B Network Details
The network architecture of the octree-based 3D U-Net we used in unconditional RGB texture generation is shown in Fig. 8. The U-Net has five hierarchical levels, corresponding to octree depths of 8, 7, 6, 5 and 4, with resolutions of $256^{3},128^{3},64^{3},32^{3},16^{3}$. The feature dimensions are set to $32,64,128,256,256$ respectively. The number of channels for input and output features is $4$ and $11$, respectively, due to the lack of material information.
@@
The network architecture of the octree-based 3D U-Net we used in text-conditioned PBR material generation is shown in Fig. 9. The U-Net has five hierarchical levels, corresponding to octree depths of 8, 7, 6, 5 and 4, with resolutions of $256^{3},128^{3},64^{3},32^{3},16^{3}$. The feature dimensions are set to $64,128,256,512,512$ respectively. The number of channels for input and output features is $4$ and $13$, respectively, as described in the main manuscript. The text feature is fed to U-Net via the octree-based multi-head cross attention mechanism. The cross attention layers are only inserted at the least two down-sampling blocks, the middle block and the two first up-sampling blocks to save GPU memory.